Left Direct Gregorian Technical Observing with IRTC: UT 2010 January 01
Happy New Year!

Observer: JHill (LBTO), DMiller (Tucson)
Telescope Operator: WWack
Instrument Support: DOfficer
Telescope Support: JUrban
SW support: (Tucson)
Phone Support: DAshby, DCox

Summary

Finally a night with good weather! Sky was mostly clear with scattered cirrus drifiting by. Seeing was suberb early with guider images as good as 0.35 arcsec FWHM. Seeing was more average 0.7 - 1.0 arcsec around the middle of the night. It was fairly stable between 0.5 and 0.7 arcsec toward the end of the night. High pressure is building so we see the ambient temperature increasing during the night.

We had some problems with the LDG rotator dropping into velocity mode during slews. Dan C. didn't have a quick fix, so he will work on it in the morning. See Issue #2475. Probably lost an hour altogether.

We had another instance of the spurious correction forces as in Issue #2443. Lost about 40 minutes of data to this.

Hardware and software worked exceedingly well. Except for the above we only lost 5 minutes to an IRAF meltdown.

We took some transformation data, and some field aberration data, and some dice5 data, and some closed-loop tracking tests.

Preparations

Ran snow melting from 11:40 to 13:40 MST to assist the sun in getting the last bit of ice off the roof.

Ran mirror ventilation to cool down toward expected night temperature, but see Issue #2474.

Doug O. moved LBC-Blue hub up 1.0 mm to a better winter focus position.

22:54 Move enclosure around to azimuth 240 to get sun on front doors, and to open side vents to cool the chamber. Telescope steel varies between -1 and +2 degC, and the present outside ambient is -2 degC, so we aren't so far off.

23:10 Ventilation mode running to get the last of the solar/snowmelt heat out of the shutters.

Details

Open

00:30 Open T=-1.5 degC, D=-38.5 degC, wind light from NW, sky clear with a very few cirrus in the far west.

00:40 Reset encoders at Az=120 El=90

00:49 pointto M5_0050 gs=0 TelM="ACQUIRE" PAM- - Rotator not ready

00:50 pointto M5_0050 gs=0 TelM="ACQUIRE" PAM- - Rotator was in velocity mode, and keeps dropping back to that. Axis servo has unexpectedly changed to velocity mode.

00:53 Wayne cycles power to rotator/chain. Problem continues. Wayne recovers with hold mode just before the preset ended.

00:57 ptspiral finds ghost of star on second step. Adjust IE=+4 CA=+76 at EL=67.2 AZ=135.4 IA=-404.

01:06 pointto M5_0050 gs=2 TelM="ACTIVE" - rotator hiccups again ERROR - Rot. disabled: State = HOLDING, TRACKING or WAIT_FOR_NEW_TARGET
Recover with rhold and rslewtotrack.

Mountain polycom unit went to bluescreen video again. We can see Doug, but he only sees blue.

01:09 pointto M5_0050 gs=2 TelM="ACTIVE" - succeeded. Seeing is 0.6 arcsec after a couple active optics updates.

Adjusting Pointing

01:14 pointto M5_0896 gs=0 TelM="ACQUIRE" PAM- ptspiral - Adjust IE=+15 CA=140 at EL=28 IA=-404

01:15 Wayne turning mirror ventilation on. Tglass_SX=-1.8 degC. Tref_SX=-1.3 degC

01:25 Adjust IA=-276 (was -404) CA=+26

01:28 pointto M5_0050 gs=0 TelM="ACQUIRE" PAM- Adjust IE=+4 CA=+27

01:33 pointto BS9105 gs=0 TelM="ACQUIRE" PAM- Adjust IE=+10 CA=+30 at EL=55 AZ=134 IA=-276

01:35 pointto BS9105 gs=6 TelM="ACTIVE" PA=INDEF=337.51

BS9107 Field Data Part I

01:43 pointto BS9107 gs=0 TelM="ACQUIRE" PAM- Adjust IE=-3 CA=+36 at EL=60.5 AZ=70 IA=-276

01:47 Clear active optics to give IDL collimation something to chew on. But note that image has only degraded to 0.75 arcsec. Thus it appears that no improvements to the default collimation model are needed.

01:48 pointto BS9107 gs=70 TelM="GUIDE"

01:51 Doug sending agw_collimate

01:54 Sent agw_collimate with and without /DONT_ASK (newly installed flag) and both work. (wfsc 1,2 and 3)

01:55 Sent agw_collimate_loop for 3 iterations (wfs 4-6) down to 200nm rms

01:58 John send pointto gs=39 TelMode="GUIDE". Rotator had to be kicked and preset sent again. IRTC star is 4" from spot, about the same as the acquisition image. There seems to be something wrong with either the new Transform, or another problem.

02:06 John runs "field_BS9107.cl". Fail to acquire image (not put in /OldRepository/AGW_Data/100101_field directory. Doug runs the agw_collimate_loop in this directory and it works fine (no permissions problems).

02:15 John called Dave A. and asks about the DG rotator always going to velocity mode on every preset where the rotator had to slew. It was failing in at least two of three slews. Dave thought it looked like a time problem (polys getting clobbered?), but time looks OK. He was going to call Dan and see if he could take a look.

02:19 Doug adds a "pwd" to agw_collimate and test. John still has troubles. Discover that John has /dontask and should be /dont_ask in agw_collimate.py

wfsc 7-10 on-axis 70 11-13 off-axis 14-15 on-axis 39 16-18 off-axis 19-20 on-axis 97 21-23 off-axis

02:40 dewpoint is down to -39 degC. Everything you do in the control room draws an arc.

Persson's Star = BS9107
Start time = 2:30

Guide Star wfscimage000*
0 7-10
70 11-13
0 14-15
39 16-18
0 19-20
97 21-23
0 24-25
82 26-28

End time = 3:02

03:04 On-axis star was off of the acquisition image. Elevation is now 75 degrees. Decide to roll back the AGW Transformation to the old (0910??) version and take some transformation data for a check. We will take another four field points aftwards (BS9107) will be on the way down in elevation in about 1.5 hours).

StoneC Transformation Data

03:10 pointto M5_0151 gs=0 ACQUIRE PAM- - Adjust IE=-1 CA=+30

03:15 pointto StoneCagw gs=0 ACQUIRE PAM-

03:20 In the process of restarting the oacserver with the "old" transformation.

what was reported from lbtmu107 in John's email of Dec 30
-rw-r--r-- 1 root root 26251 Nov 23 13:39 oacontrol.conf.20091204.agw34
-rw-r--r-- 1 root root 17793 Nov 23 13:39 oacontrol.conf.default
-rw-r--r-- 1 root root 26267 Dec  4 12:18 oacontrol.conf.agw3
-rw-r--r-- 1 root root 26267 Dec  4 17:15 oacontrol.conf.20091224
-rw-r--r-- 1 root root 26258 Dec 23 14:17 oacontrol.conf.20091223
-rw-r--r-- 1 root root 26258 Dec 23 14:17 oacontrol.conf

what we found at the start of tonight
-rw-r--r-- 1 root    root     26250 Oct 26 17:20 oacontrol.conf.StoneA
-rw-r--r-- 1 root    root     26264 Dec 16 13:28 oacontrol.conf
-rw-r--r-- 1 root    root     17793 Dec 16 13:28 oacontrol.conf.default
-rw-r--r-- 1 root    root     26255 Dec 30 16:05 oacontrol.conf.20091223

03:18 Stop GCSL, stop AGW4, stop oacserver on lbtmu107

what we change it to now
-rw-r--r-- 1 root    root     26250 Oct 26 17:20 oacontrol.conf.StoneA
-rw-r--r-- 1 root    root     26264 Dec 16 13:28 oacontrol.conf.20091216
-rw-r--r-- 1 root    root     17793 Dec 16 13:28 oacontrol.conf.default
-rw-r--r-- 1 root    root     26255 Dec 30 16:05 oacontrol.conf.20091223
-rw-r--r-- 1 root    root     26255 Dec 30 16:05 oacontrol.conf

03:30 pointto StoneCagw gs=0 ACQUIRE PAM- IE=4 CA=32

03:34 pointto StoneCagw gs=0 ACTIVE - seeing on guider is 0.35 arcsec at EL=56

03:39 pointto StoneCagw gs=20 ACTIVE

03:43 takeseq 1 40 (cube went to subdirectory of /newdata/20100101- my fault) I fixed it later.
The "Path FITS" on IRTC GUI should be simply "r:" for the cubes.

03:49 cl < home$tran_StoneC_00.cl
transtar (20)
getxy: AGW4: Network Error. (script was trying to talk to lbtmu105)

03:52 John changes the hardwired agw machine in transtar.cl to lbtmu107. This needs to become a parameter.

Significant cirrus clouds passing overhead.

Dan C. calls back, and changes our DSP config file. Restart MCSPU. This causes jet to reboot. Dan thinks there is a memory leak.

03:57 Dan restarts rpcserver, but it crashes again.

04:01 Back to another DSP configuration which stops the crashing. but we still have the velocity mode problem. We also have a problem that it doesn't alway display the LDG rotator bits on startup.

04:16 back to sky while Dan goes off to study telemetry.

04:17 pointto StoneC gs=20 ACTIVE

04:18 stop and restart MCS

04:20 pointto StoneC gs=20 ACTIVE

04:21 Stop and restart PCS, restore IE, CA, IA

04:22 pointto StoneC gs=20 ACTIVE -succeeds except for the rotator velocity mode problem.

04:30 cl < home$tran_StoneC_00.cl seeing degraded to 0.9 arcsec. We seem to have clouds.
transtar (20) irtc.20100101.043226.fits /newdata/irtc.20100101_00008_cube.fit
transtar (19) irtc.20100101.043442.fits
transtar (14) irtc.20100101.043653.fits
transtar (16) irtc.20100101.043911.fits
transtar (17) irtc.20100101.044132.fits
transtar (18) irtc.20100101.044406.fits
transtar (5) irtc.20100101.044625.fits
transtar (2) irtc.20100101.044908.fits

getdat data is in /home/agwuser/agw-control/logbook.1 on lbtmu107.

04:50 offset 30 30

04:51 takeseq 1 40 for blank sky irtc.20100101.045039.fits

04:52 Another Dan experiment that didn't work, so Dan will try to sort out the Direct rotator in the morning.

05:04 pointto

Stop and restart PCS

05:05 pointto StoneCagw gs=20 TelM="ACTIVE"

BS9107 Field Data Part II

05:08 pointto M5_0151 gs=0 TelM="ACQUIRE" PAM- - Adjust IE= 2 CA=40

05:12 pointto BS9107 gs=0 TelM="ACTIVE" - seeing is about 0.6. The thicker cirrrus clouds are gone.

05:17 pointto BS9107 gs=0 ACQUIRE PAM- Adjust IE=0 CA=29

Rearrange the order of the stars in field_BS9107.cl, although not needed as rotator angle is different now.

05:21 cl < home$field_BS9107.cl Something went wrong with rotator or preset or something, plus probe error.

ignore wfsc 29-32 the collimation got screwed up by rotator stopping

05:25 pointto BS9107 gs=0 TelM="GUIDE"

05:28 cl < home$field_BS9107.cl but interrupted to do some manual collimation. wfsc 33-36.

05:34 cl < home$field_BS9107.cl (IRAF uparm pipe error)

05:35 cl < home$field_BS9107.cl now in "Primary" mode

ERROR: cannot open `uparm$pipe10116' for writing
  "print ('!date -u "+HELPM%S"') | cl | scan (systime)"
     line 159: IRTC$pointto.cl
     called as: `pointto (sourceID=BS9107, gstarID=82, PA=27.77, TelMode=GUIDE, PAMode=yes)'

Restart IRAF window - lost 5 minutes.

05:40 cl < home$field_BS9107.cl - seeing is about 1 arcsec, but improves to 0.6 during the sequence.

Persson's Star = BS9107
Start time = 5:35

Guide Star wfscimage000*
0 37-40
0 41-44
82 45-47
0 48-49
35 50-52
0 53-54
19 55-57
0 58-59
64 60-62
0 63-64
96 65-67
0 68-69
40 70-72
0 73-74
35 75-77
0 78-79
48 80-82
0 83-84
8 85-87
0 88-89
68 90-92
0 93-94

06:22 CA=26 guessed wrong sign

06:25 CA=33

End time = 06:55

BS9122 Dice5 data

06:57 pointto BS9122 gs=0 ACQUIRE PAM-

07:02 pointto BS9122 gs=17 ACTIVE

(off1   =                   7.) Offset1 in arcsec
(off2   =                   7.) Offset2 in arcsec
(coord  =                DETXY) COORD: RADEC, AZALT, or DETXY?
(exptime=                   1.) IRTC Exposure time in seconds
(Nexpo  =                   50) Number of sequence exposures
(loopN  =                    5) Number of dice5 iterations
(verbose=                  yes) Verbose output to screen?

07:08 dice5 (as above)
position 1 seems to have failed to take an image. Position 2 is irtc.20100101.070941.fits and irtc.20100101_00017_cube.fit. Seeing is about 0.8 arcsec on guider at EL=73.

07:12 changed back to "split" mode

07:44 Loop 5 Position 5 is irtc.20100101.074332.fits and irtc.20100101_00040_cube.fit

07:45 dice5 (as above) irc OffsetPointing 0. 0. 0. DETXY MOUNT true ABS left
Loop 1 position 1 is irtc.20100101.074509.fits and irtc.20100101_00041_cube.fit

08:06 1.4 arcsec image is looking rather triangular. Was there a bad active correction? Next correction is 1687 nm rms and improves things. Astigmatism seems to be bouncing around by 1000 nm.

08:18 The problem is that we've hit a force correction limit at 07:30, so it isn't correcting the astigmatism. This is another occurrence of Issue #2443. Some of the apparently largest corrections (+-50 nm Z22) were sent around 05:30 UT while we were taking field data with IDL. Note the IDL's Z22 gain was not reduced by x4 as we did with GCS last month.

08:20 Loop 5 Position 5 is irtc.20100101.082034.fits and irtc.20100101_00065_cube.fit

08:21 Clear C00 on primary - this brings the image back down to 0.75 arcsec.

(off1   =                   7.) Offset1 in arcsec
(off2   =                   7.) Offset2 in arcsec
(coord  =                RADEC) COORD: RADEC, AZALT, or DETXY?
(exptime=                   1.) IRTC Exposure time in seconds
(Nexpo  =                   50) Number of sequence exposures
(loopN  =                    5) Number of dice5 iterations
(verbose=                  yes) Verbose output to screen?

Note that John neglected to clear the last DETXY offset before he started the RADEC dice5, so two of the positions are shifted off the IRTC.

08:24 dice5 with RADEC offsets (as above)

Loop 1 Position 1 is irtc.20100101.082530.fits and irtc.20100101_00066_cube.fit

Loop 5 Position 5 is irtc.20100101.090043.fits and irtc.20100101_00090_cube.fit

09:02 dice5 with RADEC offsets (as above)

Loop 1 Position 1 is irtc.20100101.090306.fits and irtc.20100101_00091_cube.fit

Loop 5 Position 5 is irtc.20100101.093819.fits and irtc.20100101_00115_cube.fit

StoneG Transform Data

09:40 pointto M5_0347 gs=0 TelM="ACQUIRE" PAM- - Adjust IE=19 CA=33

09:45 pointto StoneGagw gs=20 ACTIVE PA=INDEF=61.67

09:51 cl < home$tran_StoneG_00.cl - is center star too faint for IRTC? Some clouds drifting by, 1sec x 40

transtar (26) movie mode killed the IRTC data
transtar (25) irtc.20100101.095518.fits irtc.20100101_00116_cube.fit
transtar (18)
transtar (16)
transtar (19)
transtar (21)
transtar (24)
transtar (23)
transtar (20)

10:12 offset 60 60

10:13 - takeseq 1 40 - blank sky irtc.20100101.101342.fits irtc.20100101_00124_cube.fit

10:14 takeseq 2 40 - blank sky irtc.20100101.101453.fits irtc.20100101_00125_cube.fit

10:17 cl < home$tran_StoneG_30.cl - 2 sec x 40, seeing improved to 0.5 arcsec
transtar (25) irtc.20100101.101845.fits irtc.20100101_00126_cube.fit
transtar (18)
transtar (16)
transtar (19)
transtar (21)
transtar (24)
transtar (23)
transtar (20)
transtar (22) irtc.20100101.104139.fits irtc.20100101_00133_cube.fit

10:44 cl < home$tran_StoneG_90.cl - 2 sec x 40,
transtar (21) irtc.20100101.104541.fits irtc.20100101_00134_cube.fit
transtar (18)
transtar (20)
transtar (16)
transtar (22)
transtar (13)
transtar (14)
transtar (12)
transtar (10) irtc.20100101.110942.fitsirtc.20100101_00142_cube.fit

11:12 cl < home$tran_StoneG_180.cl - 2 sec x 40,
transtar (12) irtc.20100101.111446.fits irtc.20100101_00143_cube.fit
transtar (11)
transtar (10)
transtar (9)
transtar (8) bright
transtar (7) faint
transtar (6) faint - this is minimum brightness for 1 arcsec seeing
transtar (5)
transtar (4)
transtar (17) irtc.20100101.114119.fits irtc.20100101_00152_cube.fit

11:42 cl < home$tran_StoneG_270.cl - 2 sec x 40,
transtar (6) irtc.20100101.114453.fits irtc.20100101_00153_cube.fit
transtar (5)
transtar (4)
transtar (17)
transtar (1)
transtar (3)
transtar (26)
transtar (25) irtc.20100101.120539.fits irtc.20100101_00160_cube.fit

John is suspicious that the faintest stars above are sending poor active corrections and making the seeing look worse.

Closed loop tracking tests

12:08 pointto BS9152 gs=0 TelM="ACQUIRE" PAM- Hit it within 1.5 arcsec. No adjustment needed.

12:12 pointto BS9152 gs=5 TelM="ACTIVE" - star 5 is a binary - no good for transform data

12:15 takeseq 1 200 irtc.20100101.121549.fits irtc.20100101_00161_cube.fit

12:20 pointto BS9152 gs=15 TelM="ACTIVE" - 16th mag???

12:22 takeseq 1 200 irtc.20100101.122249.fits irtc.20100101_00162_cube.fit

12:27 takeseq 1 2000 irtc.20100101.122705.fits irtc.20100101_00163_cube.fit

12:35 some thicker cirrus moving in from West

13:05 pointto BS9152 gs=1 TelM="ACTIVE"

13:08 takeseq 1 2000 irtc.20100101.130829.fits irtc.20100101_00164_cube.fit
(star in different place on IRTC, so some coordinate/transform issue here)
(image quality is better, is it because guidestar is brighter?)

13:31 powered off the polycom to see if a rest helps it out.

13:33 Image quality getting worse after frame 1450 or so, but then better by 1550.

Close

13:45 Close for impending dawn. T=+2.8 degC, D=-35.5 degC, Wind 2-3 m/sec from N. Sky mostly clear. Scattered thin cirrus.

13:48 stopAGW -u 4

-- JohnHill - 01 Jan 2010
Topic attachments
I Attachment Action Size Date Who Comment
daq.20100101.jpgjpg daq.20100101.jpg manage 745 K 02 Jan 2010 - 00:03 JohnHill Plot of Telescope and Mirror Temperatures
field_BS9107.clcl field_BS9107.cl manage 3 K 16 Jan 2010 - 00:24 JohnHill IRAF script for collecting BS9107 field data
pmcc.dx.thermal.20100101.jpgjpg pmcc.dx.thermal.20100101.jpg manage 656 K 02 Jan 2010 - 00:04 JohnHill Plot of DX Primary Mirror Temperatures
pmcc.sx.thermal.20100101.jpgjpg pmcc.sx.thermal.20100101.jpg manage 513 K 02 Jan 2010 - 00:04 JohnHill Plot of SX Primary Mirror Temperatures
tran_StoneG_00.clcl tran_StoneG_00.cl manage 252 bytes 16 Jan 2010 - 00:26 JohnHill IRAF script for collecting StoneG transform data at PA=0
Topic revision: r10 - 16 Jan 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback